OpenStudy (anonymous):

If x=y^4, which expression is equal to y^16? A.x^3 B.x^4 C.x^5 D.x^12

OpenStudy (jdoe0001):

\(\bf x=y^4\implies {\color{red}{ \square }}=y?\)

OpenStudy (anonymous):

X^4

OpenStudy (jdoe0001):

\(\bf \Large { x=y^4\qquad \textit{taking }\sqrt[{\color{red}{ 4}}]{\qquad }\textit{ to both sides} \\ \quad \\ \sqrt[{\color{red}{ 4}}]{x}=\sqrt[{\color{red}{ 4}}]{y^4}\implies \sqrt[{\color{red}{ 4}}]{x}=y}\qquad {\color{olive}{ thus}} \\ \quad \\ x=y^{16}\implies {\color{red}{ \square }}=y?\)

OpenStudy (jdoe0001):

hmmm

OpenStudy (anonymous):

You are just confusing me more o.O

OpenStudy (jdoe0001):

I think I see ... what's being asked... ok

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

Ok

OpenStudy (jdoe0001):

\(\bf x=y^4\implies (x^1)^1=(y^4)^1\qquad thus \\ \quad \\ \textit{raising both sides by 4, would give us} \\ \quad \\ (x^1)^{\color{red}{ 4}}=(y^4)^{\color{red}{ 4}}\implies x^{1\cdot {\color{red}{ 4}}}=y^{4\cdot {\color{red}{ 4}}}\)

OpenStudy (anonymous):

B is the answer

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf x^4=y^{16}\)

OpenStudy (anonymous):

Thankyou